OpenStudy (angiepangie0726):

Help Pleasssee!! Fan & medal! Multiple choice What is the exact value of the exponential equation 22.1 + e5x = 100.1?

OpenStudy (angiepangie0726):

The e is e^5x

OpenStudy (photon336):

isolate e^5x first

OpenStudy (angiepangie0726):

e^5x=22.1-100.1

OpenStudy (photon336):

\[e^{5x} = 78\]

OpenStudy (angiepangie0726):

Ok, so do I ln it now??

OpenStudy (photon336):

yes take the ln of both sides and show me your work

OpenStudy (angiepangie0726):

ln e^5x=78 ln Then I know it needs to be a fraction at one point but Im not exactly sure how to extract the 5 from that power

OpenStudy (angiepangie0726):

Im pretty sure the answer is lm78/ln5but I just dont know how to get there

OpenStudy (angiepangie0726):

ln*** sorry

OpenStudy (photon336):

\[\ln(e^{5x}) = 5x | this~is~a~general~rule | \ln(e^{x}) = x \]

OpenStudy (angiepangie0726):

OHHH! Haha! I see :)

OpenStudy (photon336):

now solve for x

OpenStudy (angiepangie0726):

Hmmm, Do i make it into the fraction now? or divide

OpenStudy (angiepangie0726):

@Photon336

OpenStudy (photon336):

\[5x = \ln(78)\]

OpenStudy (photon336):

you had to take the ln of both sides

OpenStudy (angiepangie0726):

hmm ok

OpenStudy (photon336):

now what do you do?

OpenStudy (angiepangie0726):

Get the 5 on the same side??

OpenStudy (photon336):

show me

OpenStudy (angiepangie0726):

you make the 5 ln and then make it into a fraction (5 being on the bottom)

OpenStudy (photon336):

hey can you use the equation bar? \[5x = \ln(78)\] \[x = ?\]

OpenStudy (angiepangie0726):